OpenStudy (anonymous):

1. What is the quotient of x^2-16/2x^2-9x+4 / 2x^2+14x+24/4x+4 2. What is the simplified product of (x+1/x^2-25)(x+5/x^2+8x+7

OpenStudy (anonymous):

closed parthenthasie at the end

OpenStudy (whpalmer4):

You need to put in some more parentheses or use the equation editor...what you've written for the first one is \[x^2-\frac{16}{2x^2}-9x+\frac{4}{2x^2}+14x+\frac{24}{4x}+4\] and I'm positive that isn't what you actually meant.

OpenStudy (anonymous):

1. -5 x^2+11 x+4

OpenStudy (anonymous):

No?

OpenStudy (anonymous):

List the options i am not sure

OpenStudy (anonymous):

2. (x - 5)(x + 7)

OpenStudy (anonymous):

a. 1/x+3 b. 2x+2/x+3 c. 2x+2/2x^2-5x+3 d. 2(x+1)/2x^2-5x-3

OpenStudy (anonymous):

@passive13

OpenStudy (anonymous):

thats for the first one?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

D

OpenStudy (anonymous):

what about second one

OpenStudy (anonymous):

a. (x+1)/(x+5)(x+7) b.1/(x+5)(x+7) c. 1/(x-5)(x+7) d. 1/(x-5)(x-7)

OpenStudy (anonymous):

c

OpenStudy (anonymous):

is the answer correct? let me know

OpenStudy (anonymous):

ok I will